# Learning Getting-Up Policies for Real-World Humanoid Robots

### v1

@@ A-2 Rewards components in Stage II
Feet distance reward $\frac{1}{2}\Big{(}\exp(-100\left|\max(\boldsymbol{d}_{\text{feet}}-\boldsymbol{d}_{\min},-0.5)\right|)$ $+\exp(-100\left|\max(\boldsymbol{d}_{feet}-\boldsymbol{d}_{\max},0)\right|)\Big{)}$ 2
Foot orientation $\sqrt{\lVert\mathbf{g}_{xy}^{\text{feet}}\rVert}$0.5
Training Details
HumanUP contains two-stage learning, which are both trained in simulation. In Stage I, we train the discovery policy $f$ for overall 5B simulation steps in total, and 20K simulation steps for Stage II deployable policy $\pi$. All training is conducted on IsaacGym [52], and we train our policies using 4,096 paralleled environments on a single NVIDIA RTX 4090 or L40S GPU. For the getting-up task, we slow down the discovered trajectory to 8 seconds. For the rolling-over task, the trajectory is slowed down to 4 seconds. We use flat terrains in Stage I and varied terrains during Stage II training, involving flat, rough, and slopes. We follow previous works [29, 10, 77] to apply varied dynamics randomization such as friction, base CoM offset, and control delay.
Additional Qualitative Simulation Baseline Results
Fig. 7 showcases a visualization of getting up from a prone pose generated by baseline method [65]. It can be observed that this method generates motion that is highly unstable and unsafe to deploy in the real world. For example, its joints continuously jitter, the feet are stumbling and the body is keep jumping up. This indicates that this baseline [65] cannot be Sim2Real.

### v2

@@ A.2 Rewards Components in Stage II
Foot orientation $\sqrt{\lVert\mathbf{g}_{xy}^{\text{feet}}\rVert}$0.5
@@
Appendix B Additional Results
B.1 Additional Baseline Result Visualization
Fig. 7 showcases a visualization of getting up from a prone pose generated by the baseline method [70]. This method generates motion that is highly unstable and unsafe to deploy in the real world. For example, its joints continuously jitter, the feet are stumbling and the body keeps jumping up. This indicates that this baseline [70] cannot be Sim2Real.
Figure 8: Getting up from additional initial poses and against real-world external turbulence. Getting up from (a) leaning against walls, (b) sitting on the stairs. Robustness to external turbulence: (c) thumps from a hard stick, and (d) object hurling. HumanUP enables the robot to get up from initial poses other than lying, and is robust against external turbulence.
B.2 Additional Results on Getting Up From Sitting On Stairs & Leaning Against Walls
Fig. 8(a) and (b) show the simulation results of getting up from additional initial postures: sitting on the stairs and leaning against the wall. The results show that HumanUP can be generalized to more diverse initial postures in addition to lying on the ground. Besides, we find that the convergence is 4$\sim$ 4$\times$ 4 × faster than getting up from lying. We argue that getting up from sitting or leaning is easier because of the additional support from the ground, chairs, or walls.
B.3 Additional Robustness Test Against External Turbulence
Fig. 8(c) and (d) show the robustness test against external turbulence: thumps from a hard stick and object hurling. The results show that HumanUP getting-up policy is practically robust against certain external turbulence in the real world.
Appendix C Training Details
In Stage I, we train the discovery policy $f$ for overall 5B simulation steps, and 20K simulation steps for the Stage II deployable policy $\pi$. Each stage uses a regularization curriculum within, an implementation detail common to policy learning in legged locomotion literature. All training is conducted on Isaac Gym [56], and we train our policies using 4,096 paralleled environments on a single NVIDIA RTX 4090 or L40S GPU. For the getting-up task, we slow down the discovered trajectory to 8 seconds (8×8$\times$8 ×). We also tried 4×4$\times$4 × and 10×10$\times$10 ×. 4×4$\times$4 × leads to large torques and DoF velocities, and 10×10$\times$10 × does not converge. For the rolling-over task, the trajectory is slowed down to 4 seconds (selected through trials similar to the getting-up task). We use flat terrains in Stage I and varied terrains during Stage II, involving flat, rough, and slope terrains. We follow previous works [32, 10, 83] to apply varied dynamics randomization, such as base center of mass (CoM) offset and control delay.